### v1

@@
Velocity Dispersion Functions of Pressure-supported Galaxies in EAGLE Simulations with Varying AGN Feedback
@@
We investigate the stellar velocity dispersion functions (VDFs) of pressure-supported galaxies in the EAGLE cosmological simulations. The central stellar velocity dispersion is one of the fundamental dynamical tracers of the total mass of galaxy subhalos, alongside luminosity and stellar mass. Because it reflects the gravitational potential, the stellar velocity dispersion is expected to be relatively insensitive to feedback from Active Galactic Nuclei (AGN), a critical process that regulates the connection between other galaxy observables and subhalo mass. To examine the impact of AGN feedback, we analyze the VDFs from five EAGLE simulation runs, each adopting a different AGN feedback model: one ”standard”, two ”enhanced”, one ”reduced”, and one with no AGN feedback. We compute the stellar velocity dispersions of pressure-supported galaxies using member stellar particles, mimicking fiber spectroscopy. The VDFs from the standard and enhanced AGN feedback models show little difference. However, contrary to our initial expectation that the VDF shape would be largely insensitive to AGN feedback, the simulations with reduced and no AGN feedback show a significant excess of high velocity dispersion galaxies ($\sigma_{*}>200\penalty 10000\ {\rm{km\penalty 10000\ s}}^{-1}$) and a deficit of low dispersion galaxies ($100<\sigma_{*}({\rm{km\penalty 10000\ s}}^{-1})<200$), compared to those with standard or enhanced AGN feedback. The presence of high velocity dispersion galaxies in the no-AGN model arises from enhanced central star formation, due to the absence of AGN-driven gas heating or expulsion. Our results demonstrate that the shape of the theoretical VDF is sensitive to the strength of AGN feedback. These predictions offer a theoretical benchmark for future observational studies of the galaxy velocity dispersion function using large-scale spectroscopic surveys.
@@ 1 Introduction
[…]
@@
The central stellar velocity dispersion of a galaxy is one of the fundamental galaxy observables that connect the galaxy properties and the underlying DM mass. The central velocity dispersion is a kinematic measure determined by the central gravitational potential of galaxies defined by the total mass of galaxy (sub)halos including DM. Stellar velocity dispersions show good correlations with other observables including luminosity or stellar mass (e.g., Faber1976; Djorgovski1987; Cappellari2013a). More importantly, various scaling relations based on the stellar velocity dispersion indicate that stellar velocity dispersions are an excellent tracer of the total mass of galaxy subhalos. For example, the stellar velocity dispersions show a tight correlation with the dark matter velocity dispersion or the dark matter mass of subhalos (Wake2012; vanUitert2013; Bogdan2015; Zahid2016; Zahid2018; Sohn2024a). The observational studies based on strong or weak gravitational lensing also demonstrate that the stellar velocity dispersion traces the total mass of galaxy subhalos (e.g., Grillo2008; Auger2010; Utsumi2020).
@@
The stellar velocity dispersion has multiple advantages in tracing the total mass of galaxies. First of all, stellar velocity dispersion is relatively straightforward to obtain compared to other mass tracers including galaxy luminosity or stellar mass. These conventional mass tracers, galaxy luminosity or stellar mass, based on photometry are often suffering from the contamination by the line of sight objects or uncertain boundaries of galaxies (e.g., Bernardi2010; Bernardi2013; Bernardi2017). The determination of well-defined boundary is critical to measure the photometric mass tracers. However, the velocity dispersion measurement, particularly using fiber spectroscopy, only requires the obtaining spectrum at the central region of galaxies. Thus, the determination of stellar velocity dispersion is insensitive to the complicated photometric procedure. Second, measuring stellar velocity dispersion is less affected by the complicated assumptions on baryonic physics used for measuring other mass tracers. For example, stellar mass measurement has a systematic variation depending on the assumption of the slope of initial mass function (IMF), metallicity, and star formation history (e.g., Conroy2009). In contrast, these variations in underlying stellar populations have little impact on the velocity dispersion measurements (e.g., Knabel2025).
@@
One key advantage of using stellar velocity dispersion is that it provides an independent probe of complex baryonic feedback processes, particularly feedback from active galactic nuclei (AGN). AGN feedback was introduced to explain the observed deficit of massive galaxies relative to predictions from dark matter-only models of galaxy mass distributions (e.g., Ciotti1997; Silk1998; Menci2002; Benson2003; Binney2004; Silk2005; Springel2005a; Li2009). In massive galaxies, AGN feedback plays a critical role by heating or expelling cold gas through the activity of central supermassive black holes, thereby suppressing star formation (e.g., Schawinski2007; Combes2017) and limiting the growth of stellar mass. The AGN feedback reach out to the $\sim 10$ kpc and may affect the stellar mass growth within the region (e.g., Zubovas2016; Harrison2024). This feedback mechanism has been incorporated into cosmological simulations resulting in broad consistency with observed galaxy luminosity and stellar mass distributions (Schaye2015; Weigel2016; Pillepich2018). However, the detailed physical processes driving AGN feedback remain poorly understood (e.g., Wagner2012; Li2018), and its influence on stellar kinematics, particularly on the distribution of stellar velocity dispersions, has not been thoroughly explored.
@@
Here, we investigate the stellar velocity dispersion of galaxies in cosmological numerical simulations with various AGN feedback models implemented. We particularly use the stellar mass to velocity dispersion relation and the stellar velocity dispersion function (hereafter, VDF) to examine the impact of AGN feedback on the stellar velocity dispersion. The VDF is a statistical tool that displays the distribution of galaxy stellar velocity dispersions. In observations, several studies measure the VDFs for quiescent galaxies in field (Sheth2003; Choi2007; Chae2010; Montero-Dorta2017; Sohn2017b) or in galaxy clusters (Sohn2017a; Sohn2022) based on large-scale spectroscopic surveys. To provide the theoretical framework for these observed VDF measurements, Sohn2024b derive the cluster VDFs are significantly shifted toward the lower velocity dispersion compared to the observed VDFs. They suggest that understanding the origin of the difference in the observed and simulated VDFs can provide an important test for galaxy formation models implemented in cosmological simulations. We specifically examine whether variations in AGN feedback can account for the apparent offset between the observed and simulated VDFs. We also investigate how the shapes of the VDFs change across different AGN feedback models and compare them with the observed VDF to identify which models best reproduce the data.
@@
Our investigation of the role of AGN feedback on galaxy kinematics is distinctive from previous examination based on luminosity or stellar mass distribution because the simulations are not calibrated to match the stellar velocity dispersion distributions. Thus, the investigation of the stellar velocity dispersion distribution provides an independent testbed for testing AGN feedback models based on the statistical distribution of galaxy properties. Choi2007 explore the impact of AGN feedback on various galaxy properties including the velocity dispersions based on numerical simulations. Our study extends their work based on a much larger sample provided by cosmological simulations with various AGN feedback models.
@@
In Section 2, we introduce the EAGLE simulations with different AGN feedback models we used. In Section 3, we derive the velocity dispersion functions of galaxies in EAGLE simulations. Based on the comparison between VDFs, we investigate the impact of AGN feedback in the velocity and stellar mass distributions of galaxies in Section 4. We conclude our results in Section 5. Throughout the paper, we adopt the Planck cosmological parameters for EAGLE run cosmology consistent with PlanckCollaboration2014: $\Omega_{m}$ = 0.307, $\Omega_{\Lambda}$ = 0.693, $\Omega_{b}$ = 0.048, and $h$ = 0.6777 (= $H_{0}$/(100 km s-1 Mpc-1)).
@@ 2.1 EAGLE
We use Evolution and Assembly of GaLaxies and their Environments (EAGLE) cosmological simulations (Crain2015; Schaye2015; McAlpine2016) for studying the velocity dispersions of galaxies. EAGLE includes simulations covering various box sizes with various mass resolutions and baryonic physics models. Here, we use the intermediate resolution simulations with the initial baryonic particle mass $m_{\rm g}=1.81\times 10^{6}\penalty 10000\ \rm{M}_{\odot}$ and a dark matter (DM) particle mass $m_{\rm dm}=9.70\times 10^{6}\penalty 10000\ \rm{M}_{\odot}$, covering a 50 Mpc box. We use only the $z=0$ snapshot in this work. Because our primary goal is to compare the shapes of the velocity dispersion functions across different AGN feedback models, a single redshift snapshot is sufficient. Moreover, we compare the velocity dispersions of simulated galaxies with those of observed galaxies at $z<0.2$. Zahid2016 showed that the observed scaling relations between stellar velocity dispersion and stellar mass exhibit little to no redshift evolution. Therefore, using simulated galaxies from the $z=0$ snapshot provides a reasonable basis for comparison with observed galaxies at $z<0.2$.
We use simulations covering a comoving volume of (50 Mpc)3 from the suite of EAGLE simulations. We choose this volume because it includes runs implemented with various AGN feedback models (Schaye2015). RefL0050N0752 (hereafter, EAGLE-50) is the simulation that uses the standard subgrid physics and numerical techniques of the EAGLE simulation (see more details in Schaye2015). There are four other simulations with different AGN feedback, including C15AGNdT9L0050N0752 (hereafter, EAGLE-eAGN-dT), ViscHiL0050N0752 (hereafter, EAGLE-eAGN-Visc), C15AGNdT8L0050N0752 (hereafter, EAGLE-wAGN), and NoAGNL0050N0752 (hereafter, EAGLE-NoAGN).
There are two parameters that determine the AGN feedback models in EAGLE simulations: $\Delta T_{\rm AGN}$ and $C_{\rm visc}$. The five simulations we use here are based on AGN feedback models with various combinations of $\Delta T_{\rm AGN}$ and $C_{\rm visc}$. $\Delta T_{\rm AGN}$ is the temperature increase of the gas during AGN feedback. A larger $\Delta T_{\rm AGN}$ indicates a more powerful feedback event. $C_{\rm visc}$ is the parameter related to the viscosity of a subgrid accretion disk (Crain2015; Rosas-Guevara2015). The lower $C_{\rm visc}$ indicates a higher disk viscosity by definition. In a simulation with a lower $C_{\rm visc}$ model, the black hole growth and the resulting AGN feedback are delayed.
@@
Table 1: Summary of EAGLE simulations used in this study111
@@
Figure 1 displays the mass of black holes ($M_{\rm BH}$) as a function of the stellar mass of their host subhalos ($M_{*}$) in four EAGLE simulations (except for EAGLE-NoAGN). The black hole mass ($M_{\rm BH}$) is directly proportional to the feedback energy injected into the galaxies (Goubert2024). In general, $M_{\rm BH}$ increases as a function of $M_{*}$. The simulations with enhanced AGN feedback models (i.e., EAGLE-eAGN-dT and EAGLE-eAGN-Visc) show a steeper $M_{*}-M_{\rm BH}$ relation at $\log(M_{*}/\rm{M}_{\odot})>10$. In contrast, the black hole mass in EAGLE-wAGN is lower than those in the other three simulations at fixed stellar mass.
@@ 2.2 Identifying Pressure-dominated Subhalos in EAGLE
We use subhalos in EAGLE simulations identified based on the SUBFIND algorithm (Springel2001; Dolag2009). We select subhalos with a stellar mass larger than $10^{9}\penalty 10000\ \rm{M}_{\odot}$, roughly corresponding to the absolute magnitude limits of dense spectroscopic surveys for the low redshift at $0.03\leq z\leq 0.10$ (Sohn2017b; Sohn2024a). EAGLE simulations covering a (50 Mpc)3 volume we use typically contain $\sim 1700$ subhalos.
We focus on galaxies with pressure-dominated kinematics to study the central stellar velocity dispersions. In the case of rotation-dominated galaxies, the ordered rotation may significantly affect the velocity dispersion measurements (e.g., vanUitert2013). We identify pressure-dominated galaxies based on the kinematics of stellar particles of subhalos. In observations, identifying pressure- or rotation-dominated galaxies is not straightforward without spatially resolved spectroscopy. Thus, to facilitate the comparison with the observational sampling of quiescent galaxies, identification of quiescent galaxies based on low specific star formation rate (e.g., Zahid2016; Sohn2024a) in simulations is often used.
Unlike previous works, we use the kinematic properties to identify pressure-dominated galaxies because we only compare the simulated properties. Following Dubois2021, we separate the pressure- and rotation-supported galaxies based on $v/\sigma=0.5$, where $v$ and $\sigma$ indicate the rotational velocity and the velocity dispersion of stellar particles.
We derive the rotational velocity ($v$) based on the angular momentum of stellar particles. To compute $v$, we first calculate the total angular momentum vector of a galaxy ($\textbf{J}_{\rm tot}$) based on the relative positions ($\textbf{x}_{i}$) and relative velocities ($\textbf{v}_{i}$) of stellar particles with respect to the center of the galaxy:
@@
Then we derive the principal axes ($\hat{\textbf{J}}_{\rm tot}$) of the angular momentum:
@@
Here, $\hat{\textbf{J}}_{\rm tot}$ corresponds to the galaxy spin axis. Once the galaxy spin axis is defined, we then compute the angular momentum vector of each particle:
@@
The angular momentum ($\textbf{j}_{z,i}$) along with the galaxy spin axis (perpendicular to rotational plane) of each stellar particle is defined as:
@@
Additionally, we compute the projected distance on the rotational plane ($R_{i}$) of individual stellar particles from the galaxy spin axis. Based on these quantities, we finally compute rotational velocity of galaxies based on stellar particles within the stellar half-mass radius:
@@
(see also Sales2010; Sales2012).
We also compute the stellar velocity dispersion along with the galaxy spin axes we derived from Equation 4. The 1D stellar velocity dispersion is defined as:
@@
We use the stellar particles within the half-mass radius same as in the rotational velocity. Finally, we compute $v/\sigma$ of all galaxies based on the definition from Dubois2021.
@@
The $v/\sigma-M_{*}$ relations vary depending on the AGN feedback models, particularly at $M_{*}>10^{10}\penalty 10000\ \rm{M}_{\odot}$. Compared to EAGLE-50 with the normal AGN feedback model, the simulations with enhanced AGN feedback models tend to have a smaller number of rotation-dominated galaxies. In contrast, the simulations with weak or no AGN feedback contain a much larger number of rotation-dominated galaxies. The variation of the $v/\sigma$ relation depending on the AGN feedback model is consistent with the comparison in other simulations: Horizon-AGN versus Horizon-NoAGN (see Figure 3 in Dubois2016). The differences in the fractions of rotation- and pressure-dominated galaxies suggest that AGN feedback plays a role in determining galaxy kinematics. However, the variation in the $v/\sigma-M_{*}$ relation across the five EAGLE simulations is less than $\sim 10\%$ in the entire sample of $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$. Indeed, the $v/\sigma-M_{*}$ relation changes little at $M_{*}<10^{10}\penalty 10000\ \rm{M}_{\odot}$. Table 2 summarizes the number of pressure- and rotation-dominated systems in the five simulations.
Table 2: The number of total, pressure- and rotation-dominated galaxies in EAGLE simulations
@@
Figure 2: The $v/\sigma$ ratio as a function of stellar mass for subhalos in (a) EAGLE-50, (b) EAGLE-eAGN-dT, (c) EAGLE-eAGN-Visc, (d) EAGLE-wAGN, and (e) EAGLE-NoAGN. Black circles display the median $v/\sigma$ as a function of stellar mass for subhalos in each simulation. Horizon dashed lines indicate $v/\sigma=0.5$, where the rotation-dominated ($v/\sigma\geq 0.5$) and pressure-dominated ($v/\sigma<0.5$) galaxies are separated.

### v2

@@
Velocity dispersion functions of pressure-supported galaxies in EAGLE simulations with varying active galactic nucleus feedback
@@
We investigated the stellar velocity dispersion functions (VDFs) of pressure-supported galaxies in the EAGLE cosmological simulations. The central stellar velocity dispersion is one of the fundamental dynamical tracers of the total mass of galaxy subhalos, alongside luminosity and stellar mass. Because it reflects the gravitational potential, the stellar velocity dispersion is expected to be relatively insensitive to feedback from active galactic nuclei (AGNs), a critical process that regulates the connection between other galaxy observables and subhalo masses. To examine the impact of AGN feedback, we analyzed the VDFs from five EAGLE simulation runs, each adopting a different AGN feedback model: one ”standard”, two ”enhanced”, one ”reduced,” and one with no AGN feedback. We computed the stellar velocity dispersions of pressure-supported galaxies using member stellar particles, mimicking fiber spectroscopy. The VDFs from the standard and enhanced AGN feedback models show little difference. However, contrary to our initial expectation that the VDF shape would be largely insensitive to AGN feedback, the simulations with reduced and no AGN feedback show a significant excess of high velocity dispersion galaxies ($\sigma_{*}>200\penalty 10000\ {\rm{km\penalty 10000\ s}}^{-1}$) and a deficit of low velocity dispersion galaxies ($100<\sigma_{*}({\rm{km\penalty 10000\ s}}^{-1})<200$), compared to those with standard or enhanced AGN feedback. The presence of high velocity dispersion galaxies in the no-AGN model arises from enhanced central star formation, due to the absence of AGN-driven gas heating or expulsion. Our results demonstrate that the shape of the theoretical VDF is sensitive to the strength of AGN feedback. These predictions offer a theoretical benchmark for future observational studies of the galaxy VDF using large-scale spectroscopic surveys.
@@ 1 Introduction
A standard $\Lambda$ Cold Dark Matter (CDM) model suggests that the mass concentration of dark matter (DM) determines the structure formation. Within the DM halos, the gravitational interaction between DM and baryonic matter initiates the complicated formation and evolution of galaxies therein. One of the greatest challenges in studying galaxy evolution in DM halos is identifying an observable that connects the underlying DM halos and observable baryonic matter.
@@
The central stellar velocity dispersion of a galaxy is one of the fundamental galaxy observables that connect the galaxy properties and the underlying DM mass. The central velocity dispersion is a kinematic measure determined by the central gravitational potential of galaxies defined by the total mass of galaxy (sub)halos including DM. Stellar velocity dispersions show good correlations with other observables including luminosity or stellar mass (e.g., Faber and Jackson, 1976; Djorgovski and Davis, 1987; Cappellari et al., 2013; Mogotsi and Romeo, 2019; Cannarozzo et al., 2020; Ferragamo et al., 2021). More importantly, various scaling relations based on the stellar velocity dispersion indicate that stellar velocity dispersions are an excellent tracer of the total mass of galaxy subhalos. For example, the stellar velocity dispersions show a tight correlation with the DM velocity dispersion or the DM mass of subhalos (Wake et al., 2012; van Uitert et al., 2013; Bogdán and Goulding, 2015; Zahid et al., 2016, 2018; Sohn et al., 2024b). The observational studies based on strong or weak gravitational lensing also demonstrate that the stellar velocity dispersion traces the total mass of galaxy subhalos (e.g., Grillo et al., 2008; Auger et al., 2010; Utsumi et al., 2020).
@@
The stellar velocity dispersion has multiple advantages in tracing the total mass of galaxies. First of all, stellar velocity dispersion is relatively straightforward to obtain compared to other mass tracers including galaxy luminosity or stellar mass. These conventional mass tracers, galaxy luminosity or stellar mass, based on photometry are often suffering from the contamination by the line-of-sight objects or uncertain boundaries of galaxies (e.g., Bernardi et al., 2010, 2013, 2017). The determination of a well-defined boundary is critical for measuring the photometric mass tracers. However, the velocity dispersion measurement, particularly using fiber spectroscopy, only requires obtaining the spectrum at the central region of galaxies. Thus, the determination of stellar velocity dispersion is insensitive to the complicated photometric procedure. Second, measuring stellar velocity dispersion is less affected by the complicated assumptions on baryonic physics used for measuring other mass tracers. For example, stellar mass measurement has a systematic variation depending on the assumption of the initial mass function (IMF) slope, metallicity, and star formation history (e.g., Conroy et al., 2009). In contrast, these variations in underlying stellar populations have little impact on the velocity dispersion measurements (e.g., Knabel et al., 2025).
@@
One key advantage of using stellar velocity dispersion is that it provides an independent probe of complex baryonic feedback processes, particularly feedback from active galactic nuclei (AGNs). AGN feedback was introduced to explain the observed deficit of massive galaxies relative to predictions from DM-only models of galaxy-mass distributions (e.g., Ciotti and Ostriker, 1997; Silk and Rees, 1998; Menci et al., 2002; Benson et al., 2003; Binney, 2004; Silk, 2005; Springel, 2005; Li and White, 2009). In massive galaxies, AGN feedback plays a critical role by heating or expelling cold gas through the activity of central supermassive black holes, thereby suppressing star formation (e.g., Schawinski et al., 2007; Combes, 2017) and limiting the growth of stellar mass. The AGN feedback reaches out to $\sim 10$ kpc and may affect the stellar mass growth within the region (e.g., Zubovas and King, 2016; Harrison and Ramos Almeida, 2024). This feedback mechanism has been incorporated into cosmological simulations resulting in broad consistency with observed galaxy luminosity and stellar mass distributions (Schaye et al., 2015; Weigel et al., 2016; Pillepich et al., 2018). However, the detailed physical processes driving AGN feedback remain poorly understood (e.g., Wagner et al., 2012; Li et al., 2018), and its influence on stellar kinematics, particularly on the distribution of stellar velocity dispersions, has not been thoroughly explored.
@@
In this work, we investigated the stellar velocity dispersion of galaxies in cosmological numerical simulations with various AGN feedback models implemented. We particularly used the stellar mass to velocity dispersion relation and the stellar velocity dispersion function (VDF) to examine the impact of AGN feedback on the stellar velocity dispersion. The VDF is a statistical tool that displays the distribution of galaxy stellar velocity dispersions. In observations, several studies measure the VDFs for quiescent galaxies in field (Sheth et al., 2003; Choi et al., 2007; Chae, 2010; Montero-Dorta et al., 2017; Sohn et al., 2017b) or in galaxy clusters (Sohn et al., 2017a, 2022) based on large-scale spectroscopic surveys. To provide the theoretical framework for these observed VDF measurements, Sohn et al. (2024a) found that the cluster VDFs are significantly shifted toward the lower velocity dispersion compared to the observed VDFs. They suggest that understanding the origin of the difference in the observed and simulated VDFs can provide an important test for galaxy formation models implemented in cosmological simulations. We specifically examined whether variations in AGN feedback can account for the apparent offset between the observed and simulated VDFs. We also investigated how the shapes of the VDFs change across different AGN feedback models and compared them with the observed VDF to identify which models best reproduce the data.
@@
Our investigation of the role of AGN feedback in galaxy kinematics is distinctive from previous examinations based on luminosity or stellar mass distribution because the simulations were not calibrated to match the stellar velocity dispersion distributions. Thus, the investigation of the stellar velocity dispersion distribution provides an independent test bed for AGN feedback models based on the statistical distribution of galaxy properties. Choi et al. (2007) explored the impact of AGN feedback on various galaxy properties including the velocity dispersions based on numerical simulations. Our study extends their work based on a much larger sample provided by cosmological simulations with various AGN feedback models.
@@
In Section 2, we introduce the EAGLE simulations with different AGN feedback models we used. In Section 3, we derive the velocity dispersion functions of galaxies in EAGLE simulations. Based on the comparison between VDFs, we investigated the impact of AGN feedback in the velocity and stellar mass distributions of galaxies in Section 4. We conclude our results in Section 5. Throughout the paper, we adopt the Planck cosmological parameters for EAGLE-run cosmology consistent with Planck Collaboration et al. (2014): $\Omega_{m}$ = 0.307, $\Omega_{\Lambda}$ = 0.693, $\Omega_{b}$ = 0.048, and $h$ = 0.6777 (= $H_{0}$/(100 km s-1 Mpc-1)).
@@ 2.1 EAGLE
We used Evolution and Assembly of GaLaxies and their Environments (EAGLE) cosmological simulations (Crain et al., 2015; Schaye et al., 2015; McAlpine et al., 2016) to study the velocity dispersions of galaxies. EAGLE includes simulations covering various box sizes with various mass resolutions and baryonic physics models. Here, we used the intermediate-resolution simulations with the initial baryonic particle mass $m_{\rm g}=1.81\times 10^{6}\penalty 10000\ \rm{M}_{\odot}$ and DM particle mass of $m_{\rm dm}=9.70\times 10^{6}\penalty 10000\ \rm{M}_{\odot}$, covering a 50 Mpc box. We only used the $z=0$ snapshot in this work. Because our primary goal is to compare the shapes of the velocity dispersion functions across different AGN feedback models, a single redshift snapshot is sufficient. Moreover, we compare the velocity dispersions of simulated galaxies with those of observed galaxies at $z<0.2$. Zahid et al. (2016) showed that the observed scaling relations between stellar velocity dispersion and stellar mass exhibit little to no redshift evolution. Therefore, using simulated galaxies from the $z=0$ snapshot provides a reasonable basis for comparison with observed galaxies at $z<0.2$.
We used simulations covering a comoving volume of 50 Mpc3 from the suite of EAGLE simulations. We chose this volume because it includes runs implemented with various AGN feedback models (Schaye et al., 2015). RefL0050N0752 (hereafter, EAGLE-50) is the simulation that uses the standard subgrid physics and numerical techniques of the EAGLE simulation (see more details in Schaye et al., 2015). There are four other simulations with different AGN feedback, including C15AGNdT9L0050N0752 (hereafter, EAGLE-eAGN-dT), ViscHiL0050N0752 (hereafter, EAGLE-eAGN-Visc), C15AGNdT8L0050N0752 (hereafter, EAGLE-wAGN), and NoAGNL0050N0752 (hereafter, EAGLE-NoAGN).
There are two parameters that determine the AGN feedback models in EAGLE simulations: $\Delta T_{\rm AGN}$ and $C_{\rm visc}$. The five simulations we used here are based on AGN feedback models with various combinations of $\Delta T_{\rm AGN}$ and $C_{\rm visc}$. $\Delta T_{\rm AGN}$ is the temperature increase of the gas during AGN feedback. A larger $\Delta T_{\rm AGN}$ indicates a more powerful feedback event. $C_{\rm visc}$ is the parameter related to the viscosity of a subgrid accretion disk (Crain et al., 2015; Rosas-Guevara et al., 2015). The lower $C_{\rm visc}$ indicates a higher disk viscosity by definition. In a simulation with a lower $C_{\rm visc}$ model, the black-hole growth and the resulting AGN feedback are delayed.
@@
Table 1: Summary of EAGLE simulations used in this study.111
@@
Figure 1 displays the mass of black holes ($M_{\rm BH}$) as a function of the stellar mass of their host subhalos ($M_{*}$) in four EAGLE simulations (except for EAGLE-NoAGN). The black-hole mass ($M_{\rm BH}$) is directly proportional to the feedback energy injected into the galaxies (Goubert et al., 2024). In general, $M_{\rm BH}$ increases as a function of $M_{*}$. The simulations with enhanced AGN feedback models (i.e., EAGLE-eAGN-dT and EAGLE-eAGN-Visc) show a steeper $M_{*}-M_{\rm BH}$ relation at $\log(M_{*}/\rm{M}_{\odot})>10$. In contrast, the black-hole mass in EAGLE-wAGN is lower than those in the other three simulations at fixed stellar mass.
@@ 2.2 Identifying pressure-dominated subhalos in EAGLE
We used subhalos in EAGLE simulations identified based on the SUBFIND algorithm (Springel et al., 2001; Dolag et al., 2009). We select subhalos with stellar masses larger than $10^{9}\penalty 10000\ \rm{M}_{\odot}$, roughly corresponding to the absolute magnitude limits of dense spectroscopic surveys for the low redshift at $0.03\leq z\leq 0.10$ (Sohn et al., 2017b, 2024b). EAGLE simulations covering a 50 Mpc3 volume that we used typically contain $\sim 1700$ subhalos.
We focused on galaxies with pressure-dominated kinematics to study the central stellar velocity dispersions. In the case of rotation-dominated galaxies, the ordered rotation may significantly affect the velocity dispersion measurements (e.g., van Uitert et al., 2013). We identified pressure-dominated galaxies based on the kinematics of stellar particles of subhalos. In observations, identifying pressure- or rotation-dominated galaxies is not straightforward without spatially resolved spectroscopy. Thus, to facilitate the comparison with the observational sampling of quiescent galaxies, the identification of quiescent galaxies based on a low specific star formation rate (e.g., Zahid et al., 2016; Sohn et al., 2024b) in simulations is often used.
Unlike previous works, we used the kinematic properties to identify pressure-dominated galaxies because we only compared the simulated properties. Following Dubois et al. (2021), we separated the pressure- and rotation-supported galaxies based on $v/\sigma=0.5$, where $v$ and $\sigma$ indicate the rotational velocity and the velocity dispersion of stellar particles.
We derived the rotational velocity ($v$) based on the angular momentum of stellar particles. To compute $v$, we first calculated the total angular momentum vector of a galaxy ($\textbf{J}_{\rm tot}$) based on the relative positions ($\textbf{x}_{i}$) and relative velocities ($\textbf{v}_{i}$) of stellar particles with respect to the center of the galaxy:
@@
Then, we derived the principal axes ($\hat{\textbf{J}}_{\rm tot}$) of the angular momentum:
@@
Here, $\hat{\textbf{J}}_{\rm tot}$ corresponds to the galaxy spin axis. Once the galaxy spin axis is defined, we then computed the angular-momentum vector of each particle:
@@
The angular momentum ($\textbf{j}_{z,i}$) and galaxy spin axis (perpendicular to rotational plane) of each stellar particle is defined as
@@
Additionally, we computed the projected distance on the rotational plane ($R_{i}$) of individual stellar particles from the galaxy spin axis. Based on these quantities, we finally computed rotational velocity of galaxies based on stellar particles within the stellar half-mass radius:
@@
(see also Sales et al., 2010, 2012).
We also computed the stellar velocity dispersion along with the galaxy spin axes we derived from Equation 4. The 1D stellar velocity dispersion is defined as
@@
We used the stellar particles within the half-mass radius the same as in the rotational velocity. Finally, we computed the $v/\sigma$ of all galaxies based on the definition from Dubois et al. (2021).
@@
The $v/\sigma-M_{*}$ relations vary depending on the AGN feedback models, particularly at $M_{*}>10^{10}\penalty 10000\ \rm{M}_{\odot}$. Compared to EAGLE-50 with the normal AGN feedback model, the simulations with enhanced AGN feedback models tend to have a smaller number of rotation-dominated galaxies. In contrast, the simulations with weak or no AGN feedback contain a much larger number of rotation-dominated galaxies. The variation of the $v/\sigma$ relation depending on the AGN feedback model is consistent with the comparison in other simulations: Horizon-AGN versus Horizon-NoAGN (see Figure 3 in Dubois et al., 2016). The differences in the fractions of rotation- and pressure-dominated galaxies suggest that AGN feedback plays a role in determining galaxy kinematics. However, the variation in the $v/\sigma-M_{*}$ relation across the five EAGLE simulations is less than $\sim 10\%$ in the entire sample of $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$. Indeed, the $v/\sigma-M_{*}$ relation changes little at $M_{*}<10^{10}\penalty 10000\ \rm{M}_{\odot}$. Table 2 summarizes the number of pressure- and rotation-dominated systems in the five simulations.
Table 2: Number of total, pressure- and rotation-dominated galaxies in EAGLE simulations.
@@
Figure 2: $v/\sigma$ ratio as function of stellar mass for subhalos in (a) EAGLE-50, (b) EAGLE-eAGN-dT, (c) EAGLE-eAGN-Visc, (d) EAGLE-wAGN, and (e) EAGLE-NoAGN. Black circles display the median $v/\sigma$ as a function of stellar mass for subhalos in each simulation. Horizontal dashed lines indicate $v/\sigma=0.5$, where the rotation-dominated ($v/\sigma\geq 0.5$) and pressure-dominated ($v/\sigma<0.5$) galaxies are separated.